how long will it take a bus traveling at 60 miles per hour to overtake a car traveling at 40 mph if the car had a 1.5 hour head start?
Answer by jorel1380(3719)

If our bus is traveling at 60 mph, then X/60 is the amount of the distance covered by the bus before the two vehicles meet, per hour. Similarly, X/40 is the amount of the converging distance covered by the car. Therefore, we have:
X/60 + 1.5 = X/40
Multiplying through by 240, we get;
4X + 360 = 6X
360 = 2X
180 = X
Since our distance is 180 miles, the amount of time traveled by the bus is 180 miles / 60 miles/hour, or 3 hours. The bus has been traveling for 3 hours, the car has been traveling for 4.5 hours:
180=180.